OpenStudy (anonymous):

A circle has its center at the origin, and (5, -12) is a point on the circle. How long is the radius of the circle? Step by Step Please

OpenStudy (anonymous):

\[(x - h)^{2} + (y - k)^{2} = r ^{2}\]

OpenStudy (anonymous):

h and k are the points of the center. Just reverse the sign and plug them in. To get the radius, use the distance formula.

OpenStudy (anonymous):

\[\sqrt{(x _{1} - x _{2})^{2} + (y _{1} - y _{2})^{2}}\] \[\sqrt{(5 - 0)^{2} + (12 - 0)^{2}}\] \[\sqrt{(5)^{2} + (12)^{2}}\] \[\sqrt{25 + 144}\] \[\sqrt{169}\] 13 Actually, you can ignore the h and k for the equation since it's centered at the origin. \[x ^{2} + y ^{2} = r ^{2}\]

OpenStudy (anonymous):

The radius is 13.

OpenStudy (anonymous):

I had skipped a step and got confused Thanks so much!

OpenStudy (anonymous):

Sorry. I misread the question. I thought it said write the equation of a circle. lol